# Pricing energy spread options with variance gamma-driven Ornstein-Uhlenbeck dynamics

### v2

@@ 4.3 Analytical formula for cgf of the log return
The condition $\theta{\bf e}_{k}\in{\cal D}_{\bf Z}$ in Corollary 4.8 is equivalent to
$-\sqrt{\frac{\mu_{k}^{2}}{\Sigma_{kk}^{2}}+\frac{2}{\alpha_{k}\Sigma_{kk}}}-\frac{\mu_{k}}{\Sigma_{kk}}<\theta<\sqrt{\frac{\mu_{k}^{2}}{\Sigma_{kk}^{2}}+\frac{2}{\alpha_{k}\Sigma_{kk}}}-\frac{\mu_{k}}{\Sigma_{kk}}$ (4.13)
by recalling ${\cal D}_{{\bf Z}}$ is defined using the parameters under $\mathbb{P}$ in (2.5).   $\Box$
@@ 4.4.1 Forward prices
Suppose that ${\bf Z}\sim VGC^{n}_{m}({\bf b},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ or ${\bf Z}\sim WVAG^{n}(a,\boldsymbol{\alpha},\allowbreak\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ under $\mathbb{P}$, and assume ${\bf h},{\bf e}_{k}\in{\cal D}_{{\bf Z}}$. The forward price of $S_{k}$ at time $t$ with maturity time $T$ is
$F_{t}$ $=S_{k}(t)\exp({\kappa_{R_{k}(t,T)}^{\mathbb{Q}_{\bf h}}(1)}),\quad t\in[0,T],$ (4.14)
@@
$F_{t}=\mathbb{E}_{\mathbb{Q}_{{\bf h}}}[S_{k}(T){\,|\,}{\cal F}_{t}]=S_{k}(t)\mathbb{E}_{\mathbb{Q}_{{\bf h}}}[e^{R_{k}(t,T)}{\,|\,}{\cal F}_{t}].$
which gives the result if the expectation is finite, and this is provided by ${\bf e}_{k}\in{\cal D}_{{\bf Z}}$ by Proposition 4.7 or Corollary 4.8. ∎
@@
Figure 1: Forward price curve $\tau\mapsto F_{8}$ on $S_{1}$, where the MPR is ${\bf h}=\gamma(-0.1,-0.03)$, $\gamma=-1,1,10$. All parameters are otherwise taken from Section 5.1 but $b_{21}=b_{31}=0$ for all curves except the one with cyclical component.
@@ 4.4.2 Call options
Consider a call option on the univariate energy price $S_{k}$ with strike price $K$ and maturity time $T$. Suppose that ${\bf h},(\epsilon+1){\bf e}_{k}\in{\cal D}_{{\bf Z}}$, then by the Carr-Madan formula CM99 , the price of the call option at time $t$ is
@@
and $\epsilon>0$ is the damping parameter ($\alpha$ in CM99 ). Here, the cgf $\kappa^{\mathbb{Q}_{\bf h}}_{R_{k}(t,T)}(\theta)$ is obtained by Proposition 4.7 or Corollary 4.8 and its finiteness in the integrand is provided by $(\epsilon+1){\bf e}_{k}\in{\cal D}_{{\bf Z}}$ . As usual, ${\bf h}\in{\cal D}_{{\bf Z}}$ is necessary for $\mathbb{Q}_{{\bf h}}$ to exist. Provided these conditions hold, $\epsilon$ has no effect on the call price. We make further comments on the choice of $\epsilon$ in Section 5.1.
@@
Note that $(\epsilon+1){\bf e}_{k}\in{\cal D}_{{\bf Z}}$ is equivalent to $\mathbb{E}_{\mathbb{Q}_{\bf h}}[S_{k}^{\epsilon+1}]<\infty$, which is the necessary and sufficient condition for the Carr-Madan formula to be finite. For example, if the BDLP is ${\bf Z}\sim WVAG^{n}(a,\boldsymbol{\alpha},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$, then this is equivalent to
$0<\epsilon<\sqrt{\frac{\mu_{k}^{2}}{\Sigma_{kk}^{2}}+\frac{2}{\alpha_{k}\Sigma_{kk}}}-\frac{\mu_{k}}{\Sigma_{kk}}-1,$
which follows from (4.13), or alternatively, (CM99, , Section 5).  $\Box$
@@ 4.4.3 Spread options
$t\in[0,T]$, expressed as a function of the current energy price $(S_{k},S_{l})(t)$ and strike price $K$, which is possible as the LDOUP is Markov and the option is path independent. Suppose also that $\boldsymbol{\epsilon}\in{\cal D}_{{\bf Z}}$, then the Hurd-Zhou formula HZ10 gives the price with strike 1 as
@@
$\Gamma$ is the gamma function, and $\boldsymbol{\epsilon}=(\epsilon_{1},\epsilon_{2})$ are damping parameters that satisfy $\epsilon_{1}+\epsilon_{2}<-1$, $\epsilon_{2}>0$. Here, $\kappa^{\mathbb{Q}_{\bf h}}_{(R_{k},R_{l})(t,T)}(\boldsymbol{\theta})$ can be obtained by Proposition 4.5 or Corollary 4.6 by taking the $k$th and $l$th marginal component if $n\geq 3$, similar to Proposition 4.7, and its finiteness in the integrand is provided by $\boldsymbol{\epsilon}\in{\cal D}_{{\bf Z}}$. As before, provided these conditions hold, $\boldsymbol{\epsilon}$ has no effect on the spread option price. Then the spread option price for a general strike price $K$ is
@@ 5 Implementation
We now outline for a specific choice of model how the pricing formulas in Section 4.4 can be applied in an end-to-end numerical implementation that includes simulating the model, fitting the model, and pricing spread options.
5.1 Model and parameter specification
@@
The risk-free rate is $r=0.05$ and the period is $p=1$. The seasonality parameters are ${\bf b}:=(b_{01},\dots,\allowbreak b_{31},b_{02},\dots,b_{32})$, where
@@
The LDOUP parameters are $\boldsymbol{\vartheta}:=(a,\boldsymbol{\alpha},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$, where
@@
and $\boldsymbol{\eta}=-\boldsymbol{\mu}$. We refer to the seasonality and LDOUP parameters together as the model parameters. It is necessary to choose $\boldsymbol{\eta}$ such that $\mathbb{E}[{\bf Z}(1)]=\overline{\boldsymbol{\mu}}={\bf 0}$, which is equivalent to $\boldsymbol{\eta}=-\boldsymbol{\mu}$, so that the model is identifiable, otherwise both $\boldsymbol{\eta}$ and the intercept $(b_{10},b_{20})$ act as location parameters. The initial value ${\bf X}_{0}$ is an independent random vector in the stationary distribution.
The MPR is
@@
The current time is $T_{e}$, and $\bf S$} has been observed in the time interval $t\in[0,T_{e}]$ with sampling interval $\Delta$ to estimate the model parameters. At time $T_{e}$, we price spread options which have strike prices $K_{i}$, $i=1,\dots,13$, which are equally spaced values from 0.4 to 10, and maturity time $T_{o}>T_{e}$, where the time to maturity is $\tau=T_{o}-T_{e}$. We set $T_{e}=8$, $T_{o}=8.5$, $\Delta=\frac{1}{250}$ with time measured in years. The current energy price is ${\bf S}(T_{e})=(100,96)$
The Carr-Madan damping parameter for call options on $S_{1}$ and $S_{2}$ is $\epsilon=2.5$, which satisfies the conditions $(\epsilon+1){\bf e}_{1},(\epsilon+1){\bf e}_{2}\in{\cal D}_{{\bf Z}}$. The Hurd-Zhou damping parameter is $\boldsymbol{\epsilon}=(\epsilon_{1},\epsilon_{2})=(-3.5,1)$, which satisfies the conditions $\epsilon_{1}+\epsilon_{2}<-1$, $\epsilon_{2}>0$ and $\boldsymbol{\epsilon}\in{\cal D}_{{\bf Z}}$.
@@
In theory, any damping parameter satisfying the above-mentioned condition has no effect on the call and spread option prices. In practical terms, when numerically evaluating the pricing formulas, we check that changing to nearby damping parameters has virtually no effect, and any choice satisfying this will do. Further discussions on the choice of damping parameter is given in (AlSc2018, , Section 4.2).  $\Box$
5.2 Simulation
@@
5.3 Estimation
By (4.1), the seasonality parameters $(b_{0k},b_{1k},b_{2k},b_{3k})$ are estimated with linear regression on the log price $Y_{k}(t_{i})$, $i=0,1,\dots,m$, for each marginal component $k=1,2$. The estimation is valid since the errors are a zero mean stationary AR(1) process. Let the estimated seasonality parameters be $\widehat{\bf b}=(\widehat{b}_{01},\dots,\widehat{b}_{31},\widehat{b}_{02},\dots,\allowbreak\widehat{b}_{32})$ and
$\widehat{X}_{k}(t)=Y_{k}(t)-\mathopen{}\left(\widehat{b}_{0k}+\widehat{b}_{1k}t+\widehat{b}_{2k}\cos\mathopen{}\left(\frac{2\pi}{p}t\right)\mathclose{}+\widehat{b}_{3k}\sin\mathopen{}\left(\frac{2\pi}{p}t\right)\mathclose{}\right)\mathclose{},\quad k=1,2,$
@@ 5.3 Estimation
$\kappa_{{\bf Z}^{*}(\Delta)}(\boldsymbol{\theta})={\rm i}\langle(e^{\lambda\Delta}-1)\boldsymbol{\eta},\boldsymbol{\theta}\rangle+\kappa_{{\bf V}_{0}^{*}(\Delta)}(\boldsymbol{\theta})+\sum_{i=1}^{n}\kappa_{V_{k}^{*}(\Delta)}(\theta_{k}),\quad\boldsymbol{\theta}\in{\cal D}_{{\bf Z}},$
@@
In (Lu21, , Remark 9), it is noted that there was no known closed-form formula for $\Phi_{{\bf Z}^{*}(\Delta)}$ for a OU-WVAG process. We now have analytical formula and this substantially reduces the computational cost of MLE. While MLE can be used to estimate all parameters jointly by (5.6) and the analytical formula, with the 3-step procedure meant to be an approximation, it turns out the latter still preforms better in practice as it allows a 1-dimensional FFT with large $N=2^{12}$ to estimate the marginal parameters accurately at a low computational cost, while the former method may be limited to smaller $N$, such as $N=2^{8}$, since 2-dimensional FFT has a higher computational cost.   $\Box$
5.4 Calibration

### v3

@@ 4.3 Analytical formula for cgf of the log return
Define the interval
$D(b,\mu,\Sigma):=\mathopen{}\left(-\sqrt{\frac{\mu^{2}}{\Sigma^{2}}+\frac{2b}{\Sigma}}-\frac{\mu}{\Sigma},\sqrt{\frac{\mu^{2}}{\Sigma^{2}}+\frac{2b}{\Sigma}}-\frac{\mu}{\Sigma}\right)\mathclose{}.$
Using the notation in Corollary 4.3, the condition $\theta{\bf e}_{k}+{\bf h}\in{\cal D}_{\bf Z}$ in Corollary 4.8 can be written explicitly as an interval. Specifically, it is equivalent to $\theta{\bf e}_{k}\in{\cal D}_{\bf Z}^{\mathbb{Q}_{\bf h}}$, which is equivalent to
$\theta\in D(a,{[\boldsymbol{\mu}_{\bf h}]}_{k},{[{\boldsymbol{\Sigma}}_{\bf h}]}_{kk})\cap D(\beta_{k},\mu_{h_{k}},\Sigma_{h_{k}}).$
$\Box$
@@ 4.4.1 Forward prices
Suppose that ${\bf Z}\sim VGC^{n}_{m}({\bf b},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ or ${\bf Z}\sim WVAG^{n}(a,\boldsymbol{\alpha},\allowbreak\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ under $\mathbb{P}$, and assume ${\bf h},{\bf e}_{k}+{\bf h}\in{\cal D}_{{\bf Z}}$. The forward price of $S_{k}$ at time $t$ with maturity time $T$ is
$F_{k}(t,T)$ $=S_{k}(t)\exp({\kappa_{R_{k}(t,T)}^{\mathbb{Q}_{\bf h}}(1)}),\quad t\in[0,T],$ (4.13)
@@
$F_{k}(t,T)=\mathbb{E}_{\mathbb{Q}_{{\bf h}}}[S_{k}(T){\,|\,}{\cal F}_{t}]=S_{k}(t)\mathbb{E}_{\mathbb{Q}_{{\bf h}}}[e^{R_{k}(t,T)}{\,|\,}{\cal F}_{t}].$
which gives the result if the expectation is finite, and this is provided by ${\bf e}_{k}+{\bf h}\in{\cal D}_{{\bf Z}}$ by Proposition 4.7 or Corollary 4.8. ∎
@@
Figure 1: Forward price curve $\tau\mapsto F_{1}(8,8+\tau)$, where the MPR is ${\bf h}=\gamma(-0.1,-0.03)$, $\gamma=-1,1,10$. All parameters are otherwise taken from Section 5.1.1 but $b_{21}=b_{31}=0$ for all curves except the one with cyclical component.
A base load futures contract on the energy price $S_{k}$ with contract period (or delivery period) $(T_{1},T_{2}]$ consisting of $n$ days $t_{1},\dots,t_{n}\in(T_{1},T_{2}]$ and delivery price $K$ pays
$f_{T_{2}}=\frac{1}{n}\sum_{i=1}^{n}S_{k}(t_{i})-K$
at maturity time $T_{2}$. Thus, the base load futures price $\overline{F}_{k}(t,T_{1},T_{2})$ at time $t$, being the delivery price such that the contract have 0 value, is
$\overline{F}_{k}(t,T_{1},T_{2})=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\mathbb{Q}_{{\bf h}}}[S_{k}(t_{i}){\,|\,}{\cal F}_{t}],$ (4.14)
where
$\mathbb{E}_{\mathbb{Q}_{{\bf h}}}[S_{k}(t_{i}){\,|\,}{\cal F}_{t}]=\begin{cases}F_{k}(t,t_{i})&\text{if $t<t_{i}$,}\\ S_{k}(t)&\text{if $t\geq t_{i}$,}\end{cases}$ (4.15)
with the forward price $F_{k}(t,t_{i})$ given by (4.13).
@@ 4.4.2 Call options
Consider a call option on the univariate energy price $S_{k}$ with strike price $K$ and maturity time $T$. Suppose that ${\bf h},(\epsilon+1){\bf e}_{k}+{\bf h}\in{\cal D}_{{\bf Z}}$, then by the Carr-Madan formula CM99 , the price of the call option at time $t$ is
@@
and $\epsilon>0$ is the damping parameter ($\alpha$ in CM99 ). Here, the cgf $\kappa^{\mathbb{Q}_{\bf h}}_{R_{k}(t,T)}(\theta)$ is obtained by Proposition 4.7 or Corollary 4.8 and its finiteness in the integrand is provided by $(\epsilon+1){\bf e}_{k}+{\bf h}\in{\cal D}_{{\bf Z}}$ which is equivalent to $(\epsilon+1){\bf e}_{k}\in{\cal D}_{{\bf Z}}^{\mathbb{Q}_{{\bf h}}}$ . As usual, ${\bf h}\in{\cal D}_{{\bf Z}}$ is necessary for $\mathbb{Q}_{{\bf h}}$ to exist. Provided these conditions hold, $\epsilon$ has no effect on the call price. We make further comments on the choice of $\epsilon$ in Remark 5.1.
@@
Note that $(\epsilon+1){\bf e}_{k}\in{\cal D}_{{\bf Z}}^{\mathbb{Q}_{{\bf h}}}$ is equivalent to $\mathbb{E}_{\mathbb{Q}_{\bf h}}[S_{k}^{\epsilon+1}]<\infty$, which is the necessary and sufficient condition for the Carr-Madan formula to be finite. For example, if the BDLP is ${\bf Z}\sim WVAG^{n}(a,\boldsymbol{\alpha},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ under $\mathbb{P}$, then by Remark 4.9, this condition is equivalent to
$\epsilon\in(0,\infty)\cap(D(a,{[\boldsymbol{\mu}_{\bf h}]}_{k},{[{\boldsymbol{\Sigma}}_{\bf h}]}_{kk})-1)\cap(D(\beta_{k},\mu_{h_{k}},\Sigma_{h_{k}})-1).$
$\Box$
@@ 4.4.3 Spread options
$t\in[0,T]$, expressed as a function of the current energy price $(S_{k},S_{l})(t)$ and strike price $K$, which is possible as the LDOUP is Markov and the option is path independent. Suppose also that $\boldsymbol{\epsilon}+{\bf h}\in{\cal D}_{{\bf Z}}$, then the Hurd-Zhou formula HZ10 gives the price with strike price 1 as
@@
$\Gamma$ is the gamma function, and $\boldsymbol{\epsilon}=(\epsilon_{1},\epsilon_{2})$ are damping parameters that satisfy $\epsilon_{1}+\epsilon_{2}<-1$, $\epsilon_{2}>0$. Here, $\kappa^{\mathbb{Q}_{\bf h}}_{(R_{k},R_{l})(t,T)}(\boldsymbol{\theta})$ can be obtained by Proposition 4.5 or Corollary 4.6 by taking the $k$th and $l$th marginal component if $n\geq 3$, similar to Proposition 4.7, and its finiteness in the integrand is provided by $\boldsymbol{\epsilon}+{\bf h}\in{\cal D}_{{\bf Z}}$, which is equivalent to $\boldsymbol{\epsilon}\in{\cal D}_{{\bf Z}}^{\mathbb{Q}_{{\bf h}}}$. As before, provided these conditions hold, $\boldsymbol{\epsilon}$ has no effect on the spread option price. Then the spread option price for a general strike price $K$ is
@@
Remark 4.12.
For simplicity, let $n=2$. If we have an energy price process $(S^{\prime}_{1},S^{\prime}_{2})$ that is not necessarily positive, for which there exists a constant ${\bf c}\in\mathbb{R}^{2}$ that makes it positive such that $(S_{1},S_{2})(t)=(S^{\prime}_{2},S^{\prime}_{1})(t)+{\bf c}$, $t\geq 0$, satisfies the model, then there is a parity relation between spread option prices on $(S^{\prime}_{1},S^{\prime}_{2})$ and $(S_{1},S_{2})$.
Specifically, consider a spread option on $(S^{\prime}_{1},S^{\prime}_{2})$ with strike price $K^{\prime}>0$, which pays off
$f^{\prime}_{T}=(S^{\prime}_{1}(T)-S^{\prime}_{2}(T)-K^{\prime})^{+}$ (4.20)
at maturity time $T$, let $c:=c_{1}+c_{2}$, and suppose $K:=c-K^{\prime}>0$, then we have the parity relation
$(S_{1}(T)-S_{2}(T)-K)^{+}-(S^{\prime}_{1}(T)-S^{\prime}_{2}(T)-K^{\prime})^{+}=S_{1}(T)-S_{2}(T)-K.$
It follows that the spread option at time $t$ on $(S^{\prime}_{1},S^{\prime}_{2})$ with strike price $K^{\prime}$ is
$f^{\prime}_{t}((S_{1},S_{2})(t),K)=f_{t}((S_{1},S_{2})(t),K)-e^{-r(T-t)}(F_{1}(t,T)-F_{2}(t,T)-K),$ (4.21)
where $f_{t}((S_{1},S_{2})(t),K)$ is the spread option price on $(S_{1},S_{2})$ with strike price $K$ in (4.19), which can be computed using the Hurd-Zhou formula as mentioned above, and $F_{k}(t,T)$ is the forward prices of $S_{k}$, $k=1,2$, computed using (4.13). To express this in terms of the forward price of $S_{k}^{\prime}$, which is $F^{\prime}_{k}(t,T)=\mathbb{E}_{\mathbb{Q}_{{\bf h}}}[S^{\prime}_{k}(T){\,|\,}{\cal F}_{t}]$, use
$F_{k}(t,T)=F^{\prime}_{l}(t,T)+c_{k},\quad k\neq l.$ (4.22)
Also, there is no loss in generality in assuming $K>0$, a similar parity can be found if $K<0$.   $\Box$
Remark 4.13.
Throughout Section 4.4, the underlying for energy derivatives is the energy price as opposed to the forward or futures prices. However, both forms of underlying appear in the literature (for example, see bd15 ; BenthPir18 ; bsb04 ; gzlw23 for energy, bsb06 ; bs14 ; gs22 ; gss22b for forwards and futures, and benthbook ; cardur03 for both). By using (4.13), we can convert from a payoff in terms of forward or futures prices into one in terms of energy prices.   $\Box$
@@ 5 Implementation
We now outline for a specific choice of model how the pricing formulas in Section 4.4 can be applied to simulated and real data.
5.1 Simulated data example
In the simulated data example, we provide an end-to-end numerical implementation that includes simulating the model, fitting the model, and pricing spread options.
5.1.1 Model and parameter specification
We begin by specifying the true model, in which the parameters of the first and second marginal components have been chosen to approximately resemble the real data example in (benthbook, , Section 5.1) and bs14 , respectively.
@@ 5.1.1 Model and parameter specification
Let the seasonality function be
$k=1,2$, where $p=1$ is the period, and ${\bf b}:=(b_{01},\dots,\allowbreak b_{31},b_{02},\dots,b_{32})\in\mathbb{R}^{8}$ are the seasonality parameters with values
@@
The LDOUP parameters are $\boldsymbol{\vartheta}:=(a,\boldsymbol{\alpha},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ with values
@@
and $\boldsymbol{\eta}=-\boldsymbol{\mu}$. We refer to the seasonality and LDOUP parameters together as the model parameters. It is necessary to choose $\boldsymbol{\eta}$ such that $\mathbb{E}[{\bf Z}(1)]=\overline{\boldsymbol{\mu}}={\bf 0}$, which is equivalent to $\boldsymbol{\eta}=-\boldsymbol{\mu}$, so that the model is identifiable, otherwise both $\boldsymbol{\eta}$ and the intercept $(b_{01},b_{02})$ act as location parameters. The initial value ${\bf X}_{0}$ is an independent random vector in the stationary distribution.
The risk-free rate is $r=0.05$. The MPR is
@@
The current time is $T_{e}$, and $\bf S$} has been observed in the time interval $t\in[0,T_{e}]$ with sampling interval $\Delta$ to estimate the model parameters. At time $T_{e}$, we price spread options where the strike prices are $K_{i}$, $i=1,\dots,13$, which are equally spaced values from 0.4 to 10, the maturity time is $T_{o}>T_{e}$, and the time to maturity is $\tau=T_{o}-T_{e}$. We set $T_{e}=8$, $T_{o}=8.5$, $\Delta=\frac{1}{250}$ with time measured in years. The current energy price is ${\bf S}(T_{e})=(100,96)$
The Carr-Madan damping parameter for call options on $S_{1}$ and $S_{2}$ is $\epsilon=2.5$, which satisfies the conditions $(\epsilon+1){\bf e}_{1}+{\bf h},(\epsilon+1){\bf e}_{2}+{\bf h}\in{\cal D}_{{\bf Z}}$. The Hurd-Zhou damping parameter is $\boldsymbol{\epsilon}=(\epsilon_{1},\epsilon_{2})=(-3.5,1)$, which satisfies the conditions $\epsilon_{1}+\epsilon_{2}<-1$, $\epsilon_{2}>0$ and $\boldsymbol{\epsilon}+{\bf h}\in{\cal D}_{{\bf Z}}$.
@@
In theory, any damping parameter satisfying the above-mentioned condition has no effect on the call and spread option prices. In practical terms, when numerically evaluating the pricing formulas, we check that changing to nearby damping parameters has virtually no effect, and any choice satisfying this will do. If it appears no damping parameter achieves this, then for instance, the discretization parameters of the FFT method in Section 5.1.5 should be adjusted to obtain a finer grid. Additional discussions on the choice of damping parameter is given in Section 6.1 and (AlSc2018, , Section 4.2).  $\Box$
5.1.2 Simulation
@@
5.1.3 Estimation
By (4.1), the seasonality parameters $(b_{0k},b_{1k},b_{2k},b_{3k})$ are estimated with linear regression on the log price $Y_{k}(t_{i})$, $i=0,1,\dots,m$, for each marginal component $k=1,2$. The estimation is valid since the errors are a zero-mean stationary AR(1) process. Let the estimated seasonality parameters be $\widehat{\bf b}$ and the estimated seasonality function be $({\widehat{\Lambda}}_{1},{\widehat{\Lambda}}_{2})$, that is $({\Lambda}_{1},{\Lambda}_{2})$ with $^{\bf b}=\widehat{\bf b}$, and let
$\widehat{X}_{k}(t)=Y_{k}(t)-\widehat{\Lambda}_{k}(t)\quad k=1,2,$ (5.6)
@@ 5.1.3 Estimation
$\kappa_{{\bf Z}^{*}(\Delta)}(\boldsymbol{\theta})=\langle(e^{\lambda\Delta}-1)\boldsymbol{\eta},\boldsymbol{\theta}\rangle+\kappa_{{\bf V}_{0}^{*}(\Delta)}(\boldsymbol{\theta})+\sum_{i=1}^{n}\kappa_{V_{k}^{*}(\Delta)}(\theta_{k}),\quad\boldsymbol{\theta}\in{\cal D}_{{\bf Z}},$
@@
In (Lu21, , Remark 9), it is noted that there was no known closed-form formula for $\Phi_{{\bf Z}^{*}(\Delta)}$ for a OU-WVAG process. We now have an analytical formula and this substantially reduces the computational cost of MLE. While MLE can be used to estimate all parameters jointly by (5.7) and the analytical formula, with the 3-step procedure meant to be an approximation, it turns out the latter still preforms better in practice as it allows a 1-dimensional FFT with large $N=2^{12}$ to estimate the marginal parameters accurately at a low computational cost, while the former method may be limited to smaller $N$, such as $N=2^{8}$, since 2-dimensional FFT has a higher computational cost.   $\Box$
5.1.4 Calibration